please help me solve this
kate's father is three times her age, how old are kate and her father now if in 4years time the sum of their age is 68? use k to represent kates age

Answer by w_parminder(53)

Kates Present age = k
Kate's Fathers age= 3k
After 4 years of time,
Kate's age = k+4
Kate's Father's age = 3k+4
According to the question, in 4 years time the sum of their age is 68,
So, (k+4)+(3k+4) = 68
i.e. 4k+8 = 68
i.e. 4k = 68-8
i.e. 4k = 60
i.e. k = 60/4
k = 15
So
Kates Present age = 15 Yrs
Kate's Fathers age= 3*15 = 45 Yrs